OpenStudy (anonymous):

Identify the point (x1, y1) from the equation: y – 7 = 3(x + 2) (–7, 2) (7, –2) (–2,7) (2, –7)

OpenStudy (anonymous):

the slope is by definition \[m=\frac{ y1-y }{ x1-x }\] in your equation:y-7=3(x+2) you identify the slope m=3=>y-7=m(x+2)=> -7+y=m(2+x) after you multiply with (-1)=>7-y=-m(-2-x)=> THE POINT IS (7,-2)

OpenStudy (anonymous):

but cant the answer be c too

OpenStudy (anonymous):

No, and the correct answer is b as i've explained to you ^ up there |